OpenStudy (jmartinez638):

A company manufactures ball bearings for construction equipment. A machine makes the spherical bearings of radius 3 inches, and the process is accurate to 1/100th of an inch. Use differentials to produce an inequality approximating the likely volume of steel in a case containing 100 bearings.

OpenStudy (jmartinez638):

@Michele_Laino

OpenStudy (michele_laino):

Hint: total volume \(V\) is: \[\huge V = 100 \cdot \frac{{4\pi }}{3}{r^3}\] so, if we compute the first derivative, and we take the finite increments, we get: \[\huge \Delta V = 100 \cdot 4\pi {r^2} \cdot \Delta r\] where \(\Delta r= 1/100\)

OpenStudy (michele_laino):

and \(r=3\)

OpenStudy (jmartinez638):

That makes sense.